# Fast Jet Finding in Julia

### v1

@@ 1 Introduction
The current status of C++ and Python, with these languages having rather different paradigms, leads to friction and potentially awkward interfaces. Code developed in Python may not run efficiently at scale, leading to inefficient use of computing resources, then perhaps necessitating a rewrite in C++. The current generation of physicists is generally far more comfortable in Python and there is a loss of skills in C++, which is a challenge for the experiments.
@@
In this paper we report on the developments that have happened in the Julia code presented in polyglot-jets-chep23 , in particular the improvements that have resulted in the recent release of the production Julia package, JetReconstruction.jl jetreconstruction-jl-github . The package has been made more accessible to users, with comprehensive documentation and consistent interfaces. New algorithms have been introduced, specifically targeting jet reconstruction at $e^{+}e^{-}$ experiments, including reading data from Key4HEP’s EDM4hep Gaede:2022leb data format files. Support for jet substructure analysis at $pp$ colliders has been introduced. We also give the latest benchmarking results, that continue to demonstrate better performance than Fastjet Cacciari:2011ma for almost all parameters.
@@ 2 Production Release of Julia Jet Reconstruction
Details of the algorithms and strategies used for $pp$ events in the Julia version of jet finding have been described in polyglot-jets-chep23 . However, due to the different development history of how the two strategies, N2Plain and N2Tiled, were implemented, the original return values from the reconstruction were different. In the N2Plain case an implicit $p_{T}$ cut, selecting inclusive jets, was made; whereas in the N2Tiled strategy a dedicated object, called a ClusterSequence was returned. The advantage of the latter (which was inspired by Fastjet) is that it stores the entire history of the reconstruction process. Therefore it is much more useful for subsequent processing and analysis of the jet reconstruction. For the production release this was therefore unified to returning a ClusterSequence in both cases.
This choice then allowed the implementation of the other core jet selection, viz. inclusive jets. An interface was added where the same method can be used to make a selection on either the number of final jets, or on the maximum value of the metric distance ($d_{ij}$) – this takes advantage of the fact that in Julia method parameters can be named, providing a clearer interface than the type based method selection in C++.
In addition to the jet selection, another algorithm for $pp$ jet reconstruction was added, the generalised ${k}_{\text{T}}$ algorithm. This algorithm uses as a momentum metric $k^{2p}_{\text{T}}$, where the power value $p$ is arbitrary (and for specific integer powers of $-1$, 00 and $1$, maps to the well known anti-${k}_{\text{T}}$, Cambridge/Aachen, and inclusive-${k}_{\text{T}}$ algorithms, respectively).
@@
Of the two core strategies, N2Tiled scales much better to higher initial cluster densities, as found in typical LHC $pp$ events. However, there is an overhead for this tiling, which makes the N2Plain strategy better at low cluster densities. It is highly desirable that the user would not have to manually select a strategy, so a heuristic performance scan was made, indicating that the performance of each strategy is about the same for 80 input clusters. Therefore a third strategy, Best was introduced, which selects N2Plain for 80 clusters or less, otherwise N2Tiled.
Before a useful release of the software could be made, documentation for the package had to be written. This was done using the standard Julia documentation support package, Documenter.jl documenter-jl , which has the great advantage of using the inline code docstrings to document methods. Documentation was then published onto the JuliaHEP organisation’s GitHib Pages website jetreco-docs .
@@ 3.3 Implementation Details
For an optimal implementation of the $e^{+}e^{-}$ algorithms we introduce a new Julia structure to represent an jet to be reconstructed in $(\theta,\phi)$ space, an EEjet.
@@ 5 Substructure and Taggers
Table 3: Time in $\mu\mathrm{s}$ to perform various substructure and tagging operations in JetReconstruction.jl and in Fastjet (jet filtering is with a filter radius of 0.3 and selecting the 3 hardest jets; trimming is uses a trim radius and trim fraction of 0.3, with Cambridge/Aachen reclustering; mass drop uses $\mu=0.67$, $y=0.09$; and soft drop uses $z_{cut}=0.1$, $\beta=2.0$). The average is taken by performing these operations on all inclusive jets above 2GeV in all of our input sample files. Note that soft drop is not available in the Fastjet release.
@@ 6 FCCee Jets and EDM4hep
As an example of how to easily support experiment EDMs in JetReconstruction.jl we have introduced support for reading EDM4hep Gaede:2022leb events directly. Like visualisation, this support takes advantage of the Julia package manager’s concept of extensions, which means that the additional code to support this is not loaded unless the EDM4hep.jl package is present in the user’s environment. Once this is done it is trivial to call the main jet reconstruction directly using the ReconstructedParticles object from an EDM4hep ROOT file.

### v2

@@ 1 Introduction
The current status of C++ and Python, with these languages having rather different paradigms, leads to friction and potentially awkward interfaces. Code developed in Python may not run efficiently at scale, leading to inefficient use of computing resources, even necessitating a rewrite in C++. The current generation of physicists is generally far more comfortable in Python and there is a loss of skills in C++, which is a challenge for the experiments.
@@
In this paper we report on the developments that have happened in the Julia code presented in polyglot-jets-chep23 , in particular the improvements that have resulted in the recent release of the production Julia package, JetReconstruction.jl jetreconstruction-jl-github . The package has been made more accessible to users, with comprehensive documentation and consistent interfaces. New algorithms have been introduced targeting jet reconstruction at $e^{+}e^{-}$ experiments, including reading data from Key4HEP’s EDM4hep Gaede:2022leb data format files. Support for jet substructure analysis at $pp$ colliders has been introduced. We also give the latest benchmarking results, that continue to demonstrate better performance than Fastjet Cacciari:2011ma for almost all parameters.
@@ 2 Production Release of Julia Jet Reconstruction
Details of the algorithms and strategies used for $pp$ events in the Julia version of jet finding have been described in polyglot-jets-chep23 . However, due to the different development history of how the two strategies, N2Plain and N2Tiled, were implemented, the original return values from the reconstruction were different. In the N2Plain case an implicit $p_{T}$ cut, selecting inclusive jets, was made; whereas in the N2Tiled strategy a dedicated object that encapsulates the reconstruction sequence was returned that contains a record of all jet merging steps, with parents and metric distance ($d_{ij}$) parameter, allowing the complete graph of the process to be obtained. This structure, a ClusterSequence object, is much more useful as this information is required for further jet processing. This object is now used in all sequential reconstruction algorithms (but always through accessors, to hide internal implementation details).
This choice then allowed the implementation of the other core jet selection, viz. inclusive jets. An interface was added where the same method can be used to make a selection on either the number of final jets, or on the maximum value of the metric distance ($d_{ij}$) – this takes advantage of the fact that in Julia method parameters can be named, providing a clearer interface than the type based method selection in C++ (e.g., inclusive_jets(cs::ClusterSequence; ptmin=5.0) for a $p_{T}$ cut).
In addition to the jet selection, another algorithm for $pp$ jet reconstruction was added, the generalised ${k}_{\text{T}}$ algorithm. This algorithm uses as a momentum metric $k^{2p}_{\text{T}}$, where the power value $p$ is arbitrary (and for specific integer powers of $-1$, 00 and $1$, maps to the well known anti-${k}_{\text{T}}$, Cambridge/Aachen, and inclusive-${k}_{\text{T}}$ algorithms, respectively Cacciari:2005hq; Matteo_Cacciari_2008; fastjetmanual ).
@@
Of the two core strategies, N2Tiled scales better to higher initial cluster densities, as found in typical LHC $pp$ events. However, there is an overhead for this tiling, which makes the N2Plain strategy better at low cluster densities. It is highly desirable that the user would not have to manually select a strategy, so a heuristic performance scan was made, indicating that the performance of each strategy is about the same for 80 input clusters. So a third strategy, Best was introduced, which selects N2Plain for $<=80$ clusters, otherwise N2Tiled.
Before a useful release of the software could be made, documentation for the package had to be written. This was done using the standard Julia documentation support package, Documenter.jl documenter-jl , which has the great advantage of using the inline code docstrings to document methods. Documentation was then published onto the JuliaHEP organisation’s GitHub Pages website jetreco-docs .
@@ 3.3 Implementation Details
For an optimal implementation of the $e^{+}e^{-}$ algorithms we introduce a new Julia structure to represent a jet to be reconstructed in $(\theta,\phi)$ space, an EEjet.
@@ 5 Substructure and Taggers
Table 3: Time in $\mu\mathrm{s}$ to perform various substructure and tagging operations in JetReconstruction.jl and in Fastjet (jet filtering is with a filter radius of 0.3 and selecting the 3 hardest jets; trimming is uses a trim radius and trim fraction of 0.3, with Cambridge/Aachen reclustering; mass drop uses $\mu=0.67$, $y=0.09$; and soft drop uses $z_{cut}=0.1$, $\beta=2.0$). The average is taken by performing these operations on all inclusive jets above 2GeV in all of our input sample files. Note that soft drop is not available in the core Fastjet release.
@@ 6 FCCee Jets and EDM4hep
As an example of how to easily support experiment event data models (EDM) in JetReconstruction.jl we have introduced support for reading EDM4hep Gaede:2022leb events directly. Like visualisation, this support takes advantage of the Julia package manager’s concept of extensions, which means that the additional code to support this is not loaded unless the EDM4hep.jl package is present in the user’s environment. Once this is done it is trivial to call the main jet reconstruction directly using the ReconstructedParticles object from an EDM4hep ROOT file.